# TreeVQA: A Tree-Structured Execution Framework for Shot Reduction in Variational Quantum Algorithms

### v1

@@ 2.3. Energy Landscapes from Multiple VQA Tasks
A primary application of VQAs (specifically VQE) is estimating a molecule’s ground state energy - a task that scales exponentially on classical computers (Gokhale:2019) and enables predictions of molecular stability, reactivity, and dynamics. Comparing ground state energies across different molecular configurations is especially critical as it reveals insights about reaction pathways and outcomes. These relationships are captured by the potential energy surface (PES), which maps a molecule’s energy as a function of atomic positions. The PES (shown to the bottom right of Fig.3) highlights key features such as minima (stable states), maxima (transition states), and saddle points. High-accuracy ground state estimates improve the PES resolution, enabling more reliable predictions of molecular stability, vibrational spectra, and dynamic behavior (Grimsley_2019).
@@ 8.2. Fidelity Benefit for a Fixed Shot Budget
Another way to interpret TreeVQA’s benefit is through fidelity improvements under a fixed shot budget. As shown in  Figure 7, TreeVQA consistently achieves higher fidelity compared to the baseline.
Additionally, TreeVQA leads to lower fidelity variance in fidelity across different tasks (e.g., varying bond lengths or magnetic field strengths). This is because the tasks in TreeVQA are solved in a coordinated fashion, guided by shared information within clusters. In contrast, baseline VQEs, executed independently, often exhibit large disparities in fidelity under a uniform shot budget, since the optimal allocation for each task is unknown in advance.
@@ 8.4. Estimating Benefits for Large-Scale Problems
To evaluate TreeVQA on large-scale problems that may approach near-term quantum utility, we consider two benchmarks: $C_{2}H_{2}$ and a 25-site Ising model (each with ten tasks), using 28 and 50 qubits, respectively. Simulations are performed with the PauliPropagation method, truncating Pauli terms with frequencies above 8 (pauli_propagation). Results are shown in Figure 9. As classical solvers cannot provide exact ground states for these systems, we allocate a fixed number of iterations and measure the number of shots the baseline VQE requires to reach comparable energies. For some tasks, the baseline does not achieve TreeVQA’s energies within the given shots (indicated with hatch lines), suggesting that the actual shot savings will be even higher. Overall, TreeVQA demonstrates substantial shot reductions over the baseline across both benchmarks. The lower benefit observed for $C_{2}H_{2}$ compared to the Ising model partly reflects the limited computational budget allocated to $C_{2}H_{2}$ (due to runtime constraints). In fact, the average energy gap between TreeVQA and the baseline is greater for $C_{2}H_{2}$ than for the Ising model. With additional budget, the shot savings for $C_{2}H_{2}$ would likely approach those observed for the Ising. Additionally, we evaluate TreeVQA under realistic noise. We use a circuit-level noise model (Fowler_2012; caro2024learning_ptm) with error rates of $p_{\text{single-qubit}}=0.1\%$, $p_{\text{two-qubit}}=0.5\%$, and $p_{\text{readout}}=0.5\%$. For the Ising model at this scale, TreeVQA’s benefits are smaller than in the noiseless setting because neither TreeVQA nor the baseline can fully optimize in the presence of this level of noise.
@@ 8.5. TreeVQA integrated with good VQE Initialization
Although TreeVQA significantly reduces quantum hardware execution (shot) costs for VQAs, it is not the only approach to do so. Shot costs can also be reduced via classical initialization: if good VQA parameters are identified classically, quantum execution can start from these parameters, accelerating convergence. For VQE, the state-of-the-art classical initialization method is CAFQA (cafqa) and its related work (seifert2024clapton; bhattacharyya2025excitedcafqaclassicalsimulationbootstrap; bhattacharyya2023optimal), which find initial parameters using a Clifford-only search strategy. While CAFQA achieves high initialization accuracy, its benefits (and those of any classical initialization) diminish as the inherent quantum-ness of the problem increases, since classical methods can only resolve the “easy” classical component, leaving the harder quantum portion to iterative quantum execution.
@@ 9.1. Hyperparameter Analysis
We examine two key TreeVQA hyperparameters—window size and threshold—both of which will eventually affect the splitting timing. Thus we will first analyze the splitting timing effect. To further help the analysis, we introduce the Tree Critical Depth, defined as the longest path from the root to a leaf in the TreeVQA tree (see 2(b)). This metric serves as a proxy for splitting time, with more effective splitting typically resulting in shallower trees.
Analysis of Cluster Splitting Timing: Cluster splitting is crucial to TreeVQA’s performance, so we investigate how its timing impacts VQE accuracy and efficiency. We conduct experiments on three chemistry benchmarks, manually enforcing a single split during optimization; each point in Figure 13 reports the final error when this split is scheduled at different stages of optimization (expressed as a percentage of total iterations). We found that the optimal timing is application-dependent (motivating an intelligent controller) but typically occurs mid-optimization (the $H_{2}$ benchmark performs best with later splitting due to its smaller problem size, which reduces the risk of overfitting even in later stages). Splitting too early prevents subproblems from exploiting shared parameter regions, wasting resources, while splitting too late causes overfitting to the mixed Hamiltonian, requiring extra shots to escape local minima. Both cases yield fidelity losses of several percentage points under a fixed shot budget, large enough to compromise chemical accuracy.
@@ 10. Related Work
Some techniques reduce the number of Pauli measurements by grouping Pauli strings that can be measured simultaneously - either mutually qubit-wise commuting terms (gokhale2020on3) or non-commuting terms that are transformed into compatible unitary sub-Hamiltonians (yen2020measuring). ADAPT-VQE (adaptvqe) also truncates Hamiltonian terms based on their contribution to the energy, though it was originally designed for circuit-depth reduction. These approaches are orthogonal to ours and can be integrated with TreeVQA; however, they introduce their trade-offs, such as complex grouping heuristics or increased circuit depth due to additional unitary transformations.
Other work focuses on reducing the number of optimization iterations per individual task, for example, by warm-starting VQAs with classical surrogates (CAFQA_Ravi2022) or by reducing problems to simpler forms (red-qaoa). Our evaluation in Sec. 8.5 and Sec. 8.8 demonstrates that TreeVQA can seamlessly combine with these methods. However, these works do not address the multi-task structure that is inherent in real-world applications, where multiple related problems must be solved.
To our knowledge, the only prior work that explicitly targets multi-task VQE is Meta-VQE (Cervera2021). Meta-VQE designs a specialized ansatz that embeds Hamiltonian parameters directly into the circuit structure to obtain improved parameter initializations across related tasks. However, it is limited to that specific ansatz family and lacks dynamic monitoring or adaptive splitting mechanisms, meaning it cannot exploit cross-task similarity at the granularity required to achieve the substantial shot savings.

### v2

@@ 2.3. Energy Landscapes from Multiple VQA Tasks
A primary application of VQAs (specifically VQE) is estimating a molecule’s ground state energy - a task that scales exponentially on classical computers (Gokhale:2019) and enables predictions of molecular stability, reactivity, and dynamics. Comparing ground state energies across different molecular configurations is especially critical as it reveals insights about reaction pathways and outcomes. These relationships are captured by the potential energy surface (PES), which maps a molecule’s energy as a function of atomic positions. The PES (shown to the bottom right of Fig.3) highlights key features such as minima (stable states), maxima (transition states), and saddle points. High-accuracy ground state estimates improve the PES resolution, enabling more reliable predictions of molecular stability, vibrational spectra, and dynamic behavior (adaptvqe).
@@ 8.2. Fidelity Benefit for a Fixed Shot Budget
Another way to interpret TreeVQA’s benefit is through fidelity improvements under a fixed shot budget. As shown in  Figure 7, TreeVQA consistently achieves higher fidelity compared to the baseline. Additionally, TreeVQA leads to lower fidelity variance in fidelity across different tasks (e.g., varying bond lengths or magnetic field strengths). This is because the tasks in TreeVQA are solved in a coordinated fashion, guided by shared information within clusters. In contrast, baseline VQEs, executed independently, often exhibit large disparities in fidelity under a uniform shot budget, since the optimal allocation for each task is unknown in advance.
@@ 8.4. Estimating Benefits for Large-Scale Problems
To evaluate TreeVQA on large-scale problems that may approach near-term quantum utility, we consider two benchmarks: $C_{2}H_{2}$ and a 25-site Ising model (each with ten tasks), using 28 and 50 qubits, respectively. Simulations are performed with the PauliPropagation method, truncating Pauli terms with frequencies above 8 (pauli_propagation). Results are shown in Figure 9. As classical solvers cannot provide exact ground states for these systems, we allocate a fixed number of iterations and measure the number of shots the baseline VQE requires to reach comparable energies. For some tasks, the baseline does not achieve TreeVQA’s energies within the given shots (indicated with hatch lines), suggesting that the actual shot savings will be even higher. Overall, TreeVQA demonstrates substantial shot reductions over the baseline across both benchmarks. The lower benefit observed for $C_{2}H_{2}$ compared to the Ising model partly reflects the limited computational budget allocated to $C_{2}H_{2}$ (due to runtime constraints). In fact, the average energy gap between TreeVQA and the baseline is greater for $C_{2}H_{2}$ than for the Ising model. With additional budget, the shot savings for $C_{2}H_{2}$ would likely approach those observed for the Ising. Additionally, we evaluate TreeVQA under a noise model similar to that of (rudolph2025paulipropagation_error_mitigation_notebook), in which a depolarizing noise layer (with error rates of $1\%$) is inserted after each circuit repetition. Under this noise model, the advantages of TreeVQA are marginally reduced relative to the noiseless case, yet it continues to deliver substantial shot savings compared to the baseline. Large-scale noise deforms the VQA optimization landscape and introduces additional local minima, adversely affecting both TreeVQA and the baseline and limiting their ability to converge to the ground state. This increased heterogeneity in the landscape slightly constrains TreeVQA’s ability to fully exploit task similarity. As quantum hardware advances in the EFT regime and noise levels continue to decrease, we expect TreeVQA’s performance gains to progressively approach those observed in noiseless simulations.
@@ 8.5. TreeVQA integrated with good VQE Initialization
Although TreeVQA significantly reduces quantum hardware execution (shot) costs for VQAs, it is not the only approach to do so. Shot costs can also be reduced via classical initialization: if good VQA parameters are identified classically, quantum execution can start from these parameters, accelerating convergence. For VQE, the state-of-the-art classical initialization method is CAFQA (CAFQA_Ravi2022) and its related work (seifert2024clapton; bhattacharyya2025excitedcafqaclassicalsimulationbootstrap; bhattacharyya2023optimal), which find initial parameters using a Clifford-only search strategy. While CAFQA achieves high initialization accuracy, its benefits (and those of any classical initialization) diminish as the inherent quantum-ness of the problem increases, since classical methods can only resolve the “easy” classical component, leaving the harder quantum portion to iterative quantum execution.
@@ 9.1. Hyperparameter Analysis
We examine two key TreeVQA hyperparameters—sliding window size and threshold—both of which will eventually affect the splitting timing. Thus we will first analyze the splitting timing effect. To further help the analysis, we introduce the Tree Critical Depth, defined as the longest path from the root to a leaf in the TreeVQA tree (see 2(b)). This metric serves as a proxy for splitting time, with more effective splitting typically resulting in shallower trees.
Analysis of Cluster Splitting Timing: Cluster splitting timing is crucial to TreeVQA’s performance, so we investigate how it impacts VQE accuracy and efficiency. We conduct experiments on three chemistry benchmarks, manually enforcing a single split during optimization; each point in Figure 13 reports the final error when this split is scheduled at different stages of optimization (expressed as a percentage of total iterations). We found that the optimal timing is application-dependent (motivating an intelligent controller) but typically occurs mid-optimization (the $H_{2}$ benchmark performs best with later splitting due to its smaller problem size, which reduces the risk of overfitting even in later stages). Splitting too early prevents subproblems from exploiting shared parameter regions, wasting resources, while splitting too late causes overfitting to the mixed Hamiltonian, requiring extra shots to escape local minima. Both cases yield fidelity losses of several percentage points under a fixed shot budget, large enough to compromise chemical accuracy.
@@ 10. Related Work
Measurement reduction techniques often group Pauli strings that are either mutually qubit-wise commuting (gokhale2020on3) or transformed into compatible unitary sub-Hamiltonians (yen2020measuring). Others, like ADAPT-VQE (adaptvqe), truncate terms based on energy contributions. While both orthogonal and compatible with TreeVQA, these approaches introduce trade-offs, such as complex grouping heuristics or increased circuit depth from additional unitaries. Similarly, iteration reduction methods—using classical surrogates for warm-starts (CAFQA_Ravi2022) or problem simplification (red-qaoa)—seamlessly integrate with our work (Sec. 8.58.8) but fail to address inherent multi-task structures. To our knowledge, Meta-VQE (Cervera2021) remains the only prior work explicitly targeting multi-task VQE by embedding Hamiltonians parameters directly into a specialized ansatz, it lacks the dynamic monitoring and adaptive splitting mechanisms required to achieve the substantial shot savings offered by TreeVQA.